A shelter Island ferry boat moves with a cosntant velocity vox =8 m/sfor 60 sec.It then shuts off its engines and coasts.Its coasting velocity is given by vx = vox t21 /t2,

where t1= 60 s.What is the displacement of the boat for the interval 0<t<∞?


1
Expert's answer
2021-10-10T16:04:46-0400

Displacement in 60 s:


"\Delta x_1=v_{0x}T^2/T=480\text{ m}."

Displacement during the other interval of time:


"\Delta x_2=\int^\infty_Tv_x\text dt=\\\space\\
=\int^\infty_T\frac{v_{0x}T^2}{t^2}\text dt=\\\space\\
=v_{0x}T^2\int^\infty_T\frac{1}{t^2}\text dt=\\\space\\
v_{0x}T^2(-1/t)|^\infty_T=480\text{ m}."

The total displacement is


"\Delta x=\Delta x_1+\Delta x_2=960\text{ m}."